# Investigating Subtler Biases in LLMs: Ageism, Beauty, Institutional, and Nationality Bias in Generative Models

### v2

@@ Abstract
LLMs are increasingly powerful and widely used to assist users in a variety of tasks. This use risks the introduction of LLM biases to consequential decisions such as job hiring, human performance evaluation, and criminal sentencing. Bias in NLP systems along the lines of gender and ethnicity has been widely studied, especially for specific stereotypes (e.g., Asians are good at math). In this paper, we investigate bias along less-studied but still consequential, dimensions, such as age and beauty, measuring subtler correlated decisions that LLMs make between social groups and unrelated positive and negative attributes. We ask whether LLMs hold wide-reaching biases of positive or negative sentiment for specific social groups similar to the “what is beautiful is good” bias found in people in experimental psychology Dion et al. (1972). We introduce a template-generated dataset of sentence completion tasks that asks the model to select the most appropriate attribute to complete an evaluative statement about a person described as a member of a specific social group. We also reverse the completion task to select the social group based on an attribute. We report the correlations that we find for 4 cutting-edge LLMs. This dataset can be used as a benchmark to evaluate progress in more generalized biases and the templating technique can be used to expand the benchmark with minimal additional human annotation. [^1] Our dataset and code will be available at https://github.com/kamruzzaman15/Identifying-Subtler-Biases-in-LLMs.
@@
Mahammed Kamruzzaman University of South Florida kamruzzaman1@usf.edu
Md. Minul Islam Shovon Rajshahi University of Engineering and Technology mainulislam588@gmail.com                        Gene Louis Kim University of South Florida genekim@usf.edu
@@ 1 Introduction
Social scientists found that human biases extend beyond simple stereotypes and can lead to general associations of positive attributes to members holding (or perceived to hold) certain key characteristics Dion et al. (1972); Gross and Crofton (1977); Butler (1969). For example, Dion et al. (1972) found that people are more likely to infer a plethora of other desirable characteristics to people that are judged more attractive.
In this paper, we extend the evaluation of bias in LLMs in the following key ways: we investigate whether LLMs make general associations between stereotyped categories and unrelated positive, negative, and neutral attributes rather than specific stereotype inferences. In addition, we investigate dimensions of bias that have been largely overlooked: age, beauty, academic institution, and nationality. While nationality and ageism bias have been studied in a few papers, this particular setting of investigation hasn’t been explored extensively, so we can refer to these as overlooked domains together. Figure 1 shows an example of how we formulate the completion task for LLMs in the bias domain of academic institutions.
@@ 2 Motivating Studies from Social Science
Dion et al. (1972) found that people who are considered attractive are more likely to believed to hold socially desirable traits and higher occupational status. Gross and Crofton (1977) found that such biases also occur in the opposite direction—that knowing that people have desirable qualities lead us to judge them as more beautiful. Attractiveness plays a measurable role in employee termination Commisso and Finkelstein (2012), hiring processes Marlowe et al. (1996), interview callbacks Maurer-Fazio and Lei (2015), and even in victim blaming Weber et al. (2013).
Perdue and Gurtman (1990), Axelrod and Eisdorfer (1961), and Aaronson (1966) found that people are more likely to infer negative traits and stereotypes on people that are older or labeled as “old” as opposed to “young”. This validates Butler’s (1969) earlier introduction of the term “ageism”, referring to discrimination towards elderly people. As with beauty, age has been found to be a factor in evaluating workers Ng and Feldman (2012).
@@
As LLMs are trained on data created by humans, we hypothesize that they are prone to similar biases to those identified in people. If present, we must measure the degree to which such biases are present so that we can appropriately account for them when LLMs are used for consequential decisions, such as recruiting and hiring.
@@ 4 Task Definition
We consider a system to be biased if the conditional likelihood of positive, negative, or neutral completions change when the polarity of the prompting text changes. That is, an unbiased system will have $\Delta$PL, $\Delta$NL, and $\Delta$NuL values (defined in Figure 3) of 0. For example, in Figure 3, we find that $\Delta$NL is systematically negative, so the negative predictions of the models are biased in the opposite direction of the prompt text changes.
@@ 6 Experimental Setup
We use GPT-4, PaLM-2, Llama-2-13B, and Mistral-7B in our experiments. [^16] We use 4-bit quantized versions for Llama-2 and Mistral 7B due to our resource constraints., [^17] We also have experimental results for GPT-3.5, which follow the same broad trends as the other models discussed here. Results for GPT-3.5 are only in the appendices due to space constraints. For model details, see Appendix B. Exact experimental details are in our supplementary materials, including scripts for replication. We examine how LLMs respond to positive and negative attributes and stimuli. Specifically, we calculate the conditional likelihood of a model selecting positive, negative, and neutral attributes in response to stereotypically positive and negative stimuli. We will refer to these as [stimulus]-to-[attribute] likelihoods. For example, we call the likelihood of the model to select positive attributes (e.g., friendly, motivated, creative, etc.) in response to stereotypically negative stimuli (e.g., unattractive, Hinds Community College, South Sudan, 65 years old, etc.) the negative-to-positive likelihood (NPL). Our shorthand uses P for positive, N for negative, and Nu for neutral.
@@ 7.1 Addressing Possible Confounds
A key possible confounding variable in our investigation of institutional bias is educational level. We see this in Figure 4 which shows the percentage of selecting positive, negative, and neutral attributes/stimuli averaged across all models for different educational levels. In the SAI direction, models select considerably more positive and fewer negative attributes for teachers compared to first and second-year students. This correlation is also statistically significant for most of the models, see Table 28 in Appendix E.
@@
Another possible confound in our experiments is gender bias. Figure 5 shows the percentage of selecting positive, negative, and neutral attributes/stimuli averaged across all models for gendered pronoun settings. In both SAI and ASA directions LLMs are marginally more likely to select positive attributes and less likely to select negative attributes for feminine pronouns compared to masculine or non-binary pronouns. This correlation is also statistically significant for most of the models, see Table 29 in Appendix E.
@@ Appendix C Detailed Result Section
We represent our complete experimental results for each bias type and direction in Tables 8 -27.
@@ C.1 Tables in SAI direction
In Table 23, 24, 25, 26, 27 we present our results for different educational level for GPT-3.5, GPT-4, PaLM-2, Llama-2, and Mistral models respectively.
@@ Appendix D Invalid Results Categorized
Appendix E Statistical Tests for Educational Level and Gender
Model Direction $\tau$ $p$ $H_{0}$?
GPT-4 SAI 0.140 7.5e-15 Reject
ASA -0.115 1.7e-11 Reject
PaLM-2 SAI 0.132 2.6e-12 Reject
ASA 0.003 0.8582 Reject Fail
Llama-2 SAI 0.058 0.0025 Reject
ASA 0.008 0.9692 Reject Fail
Mistral SAI 0.041 0.024 Reject
ASA -0.062 0.004 Reject
Model Direction $\tau$ $p$ $H_{0}$?
GPT-4 SAI 0.031 0.0020 Reject
ASA 0.023 0.0468 Reject
PaLM-2 SAI 0.025 0.0228 Reject
ASA 0.016 0.1758 Reject Fail
Llama-2 SAI 0.026 0.0160 Reject
ASA 0.007 0.5380 Reject Fail
Mistral SAI 0.011 0.3076 Reject Fail
ASA 0.027 0.0206 Reject
Table 28 shows that the correlation between education level and attribute quality (SAI direction) is statistically significant for all 4 models. This pattern does not generally hold in the ASA direction, and in fact, there are small effect negative correlations for GPT 4 and Mistral predictions. We would not expect the education level to be predictive of the type of institution since they all have first-year students, second-year students, and teachers.
Table 29 partially confirms this observation with statistically significant results for GPT-4 in both inference directions and the other models in one of the directions. We find that the effect sizes are small even where results are significant. [^23] This is promising regarding the progress of the field, but this is only a coarse-grained analysis. We only use pronouns to access this variable and our dataset does not focus on gender bias-specific attributes and stimuli.

### v3

@@ Abstract
LLMs are increasingly powerful and widely used to assist users in a variety of tasks. This use risks introducing LLM biases into consequential decisions such as job hiring, human performance evaluation, and criminal sentencing. Bias in NLP systems along the lines of gender and ethnicity has been widely studied, especially for specific stereotypes (e.g., Asians are good at math). In this paper, we investigate bias along less-studied but still consequential, dimensions, such as age and beauty, measuring subtler correlated decisions that LLMs make between social groups and unrelated positive and negative attributes. Although these subtler biases are understudied they follow people as much as gender and ethnicity do. So, we want to see whether they also follow one with LLMs. We introduce a template-generated dataset of sentence completion tasks that asks the model to select the most appropriate attribute to complete an evaluative statement about a person described as a member of a specific social group. We also reverse the completion task to select the social group based on an attribute. We report the correlations that we find for 4 cutting-edge LLMs. This dataset can be used as a benchmark to evaluate progress in more generalized biases and the templating technique can be used to expand the benchmark with minimal additional human annotation.  [^1] Our dataset and code is available at https://github.com/kamruzzaman15/Identifying-Subtler-Biases-in-LLMs. 00footnotetext: This work has been accepted at ACL 2024.
@@ 1 Introduction
Social scientists found that human biases extend beyond simple stereotypes and can lead to general associations of positive attributes to members holding (or perceived to hold) certain key characteristics. For example, Dion et al. (1972) found that people are more likely to infer a plethora of other desirable characteristics to people that are judged more attractive—a result that has been confirmed and elaborated upon for the present day by more recent studies Commisso and Finkelstein (2012); Peng et al. (2019); Maurer-Fazio and Lei (2015); Weber et al. (2013).
In this paper, we extend the evaluation of bias in LLMs in the following key ways: we investigate whether LLMs make general associations between stereotyped categories and unrelated positive, negative, and neutral attributes rather than specific stereotype inferences. In addition, we investigate dimensions of bias that have been largely overlooked: age, beauty, academic institution, and nationality. Although understudied, these dimensions of bias follow people as much as gender and ethnicity do. Furthermore, the few existing studies regarding these biases in LLMs all study them in terms of specific stereotypes. Figure 1 shows an example of how we formulate the completion task for LLMs in the bias domain of academic institutions.
@@ 2 Motivating Studies from Social Science
Many studies demonstrate the pervasive impact of social biases in various spheres of life. Dion et al. (1972) found that people who are considered attractive are more likely to be believed to hold socially desirable traits and higher occupational status. This attractiveness bias also occurs inversely, as shown by Gross and Crofton (1977), where known desirable qualities influence our perceptions of an individual’s beauty. These two papers inspired us to set up our experimental design. More recent psychological research has confirm using modern scientific standards that such biases extend in specific real-world settings like employee termination Commisso and Finkelstein (2012), hiring practices Peng et al. (2019), interview callbacks Maurer-Fazio and Lei (2015), and even in victim blaming scenarios Weber et al. (2013).
An opposite bias effect for beauty exists for people in particular contexts, in which a highly attractive person is dispreferred Agthe et al. (2010). These effects are subject to interactions with sexual motive, social competition, culture, and the nature of the interaction making them more difficult to study Wan and Wyer (2015); Li et al. (2020). Generally, this opposing beauty bias effect appears when a preference for beauty would introduce a social threat (e.g., hiring a same-sex person, an embarrassing purchase from an opposite-sex person, etc.). In this paper we focus on the most general beauty bias where beauty is unduly associated with positive attributes leaving the opposite bias effect for future work.
Age-related stereotypes similarly impact societal perceptions and actions, with studies by Perdue and Gurtman (1990), Marques et al. (2020), and Donizzetti (2019) indicating a general propensity to infer negative traits based on age. This aligns with the foundational work on ageism by Butler (1969) and has been found to have implications in professional settings Ng and Feldman (2012).
@@
As LLMs are trained on data created by humans, we hypothesize that they are prone to similar biases to those identified in people. If present, we must measure the degree to which such biases are present so that we can appropriately account for them when LLMs are used for consequential decisions, such as recruiting and hiring. In order to do this, we measure our bias in terms of “representational harms” (using the terminology from Blodgett et al. (2020)), that is, an LLM is biased if it makes general associations between stereotyped categories and unrelated positive, negative, and neutral attributes. Following the guidance of Blodgett et al. (2020), our objective is to define and measure these biases, particularly how they manifest and potentially perpetuate existing social hierarchies.
@@ 6 Experimental Setup
We use GPT-4, PaLM-2, Llama-2-13B, and Mistral-7B in our experiments. [^16] We use 4-bit quantized versions for Llama-2 and Mistral 7B due to our resource constraints., [^17] We also have experimental results for GPT-3.5, which follow the same broad trends as the other models discussed here. Results for GPT-3.5 are only in the appendices due to space constraints. For model details, see Appendix B. Exact experimental details are in our supplementary materials, including scripts for replication. We examine how LLMs respond to positive and negative attributes and stimuli. Specifically, we calculate the conditional likelihood of a model selecting positive, negative, and neutral attributes in response to stereotypically positive and negative stimuli. We will refer to these as [stimulus]-to-[attribute] likelihoods. For example, we call the likelihood of the model to select positive attributes (e.g., friendly, motivated, creative, etc.) in response to stereotypically negative stimuli (e.g., unattractive, Hinds Community College, South Sudan, 65 years old, etc.) the negative-to-positive likelihood (NPL). Our shorthand uses P for positive, N for negative, and Nu for neutral. We consider a system to be biased if the conditional likelihood of positive, negative, or neutral completions change when the polarity of the prompting text changes. That is, an unbiased system will have $\Delta$PL, $\Delta$NL, and $\Delta$NuL values (defined in Figure 3) of 0. For example, in Figure 3, we find that $\Delta$NL is systematically negative, so the negative predictions of the models are biased in the opposite direction of the prompt text changes.
@@ 7.1 Addressing Possible Confounds
A key possible confounding variable in our investigation of institutional bias is educational level. We see this in Figure 4 which shows the percentage of selecting positive, negative, and neutral attributes/stimuli averaged across all models for different educational levels. In the SAI direction, models select considerably more positive and fewer negative attributes for teachers compared to first and second-year students. So, there is a trend toward positive representation for teacher. Table 5 shows that the correlation between education level and attribute quality (SAI direction) is statistically significant for all 4 models. This pattern does not generally hold in the ASA direction, and in fact, there are small effect negative correlations for GPT 4 and Mistral predictions. We would not expect the education level to be predictive of the type of institution since they all have first-year students, second-year students, and teachers.
Model Direction $\tau$ $p$ $H_{0}$?
GPT-4 SAI 0.140 7.5e-15 Reject
ASA -0.115 1.7e-11 Reject
PaLM-2 SAI 0.132 2.6e-12 Reject
ASA 0.003 0.8582 Reject Fail
Llama-2 SAI 0.058 0.0025 Reject
ASA 0.008 0.9692 Reject Fail
Mistral SAI 0.041 0.024 Reject
ASA -0.062 0.004 Reject
@@
As most of the institutions contain distinct words “University" and “Community College", one of the reviewers wonders if those two words dominate the selection of attributes, instead of the actual university and community college. To address this we perform a follow-up experiment where we analyzed the results for positive institutions that do not include the word “university” (e.g., MIT) and negative institutions that do not include the phrase “community college” (e.g., Dallas College). The overall trend is consistent with that of Table 2’s institutional bias, with the only difference being that in this experimental setup, in the SAI direction, the Mistral model rejects the null hypothesis, and there is no statistically significant result for that. See Table 31 in Appendix E for full results.
@@
Another possible confound in our experiments is gender bias. Figure 5 shows the percentage of selecting positive, negative, and neutral attributes/stimuli averaged across all models for gendered pronoun settings. In both SAI and ASA directions LLMs are marginally more likely to select positive attributes and less likely to select negative attributes for feminine pronouns compared to masculine or non-binary pronouns. Table 6 partially confirms this observation with statistically significant results for GPT-4 in both inference directions and the other models in one of the directions. We find that the effect sizes are small even where results are significant. [^20] This is promising regarding the progress of the field, but this is only a coarse-grained analysis. We only use pronouns to access this variable and our dataset does not focus on gender bias-specific attributes and stimuli.
@@
Model Direction $\tau$ $p$ $H_{0}$?
GPT-4 SAI 0.031 0.0020 Reject
ASA 0.023 0.0468 Reject
PaLM-2 SAI 0.025 0.0228 Reject
ASA 0.016 0.1758 Reject Fail
Llama-2 SAI 0.026 0.0160 Reject
ASA 0.007 0.5380 Reject Fail
Mistral SAI 0.011 0.3076 Reject Fail
ASA 0.027 0.0206 Reject
@@ 9 Limitations
Acknowledgements
This project was fully supported by the University of South Florida. We thank the reviewers for their valuable feedback especially their suggestions to add different types of confounding variables to our experiments.
@@ Appendix C Detailed Result Section
We represent our complete experimental results for each bias type and direction in Tables 10 -29.
@@ C.1 Tables in SAI direction
In Table 25, 26, 27, 28, 29 we present our results for different educational level for GPT-3.5, GPT-4, PaLM-2, Llama-2, and Mistral models respectively.
@@ Appendix D Invalid Results Categorized
Appendix E Statistical Tests for Educational Level
@@
GPT-4 PaLM-2 Llama-2 Mistral
DOE BT $\tau$ $p$ $H_{0}$? $\tau$ $p$ $H_{0}$? $\tau$ $p$ $H_{0}$? $\tau$ $p$ $H_{0}$?
SAI I 0.165 8.0e-08 R 0.375 3.8e-13 R 0.321 4.2e-15 R 0.038 0.2995 RF
ASA I 0.270 1.6e-21 R 0.408 7.3e-24 R 0.307 3.7e-24 R 0.105 0.0004 R
Table 31: Kendall’s $\tau$ test results for Institutional Bias when positive institutions don’t include word university and negative institutions don’t include the phrase community college. We use a significance level of $\alpha<0.05$ to reject the null hypothesis. Here, BT stands for Bias Type, I stands for Institution, R stands for Reject and RF stands for Reject Fail.